Blue arm optical alignment

Loading Map....

Date/Time
Date(s) - 06/12/2016 - 08/12/2016
00:00

Location
Geneva Observatory

Categories


Merate colleagues were in Geneva for a system integration mission:

Attendees

Marco Riva
Manuele Moschetti
Ian Hughes
Francesco Pepe

Day 1

Laser metrology and realignment
Main collimator adjustment to the nominal position (1.5 arcmin along vertical direction)
Getting confidence with the CCD and full acquisition system
Preliminary test with flat field
Order finding through green (520) and blue (460), preliminary results is 0.5 of fsr shift

Day 2

APSU reposition and reacquisition to be in nominal position (Screws were too loose like in the BTM and MC now more tight)
ThAR acquisition and order finding through harps similarity (Francesco, Danuta and Christophe)
Further test with flat field strange shadow appeared
green laser mounted in the front end and check of the footprint: laterally is good (within+-2mm) vertically ssems ok but near to the prism top edge
FE1 filter wheel dismount and notice that the strange trigger could be due to a manufacturing issue. Dismounted to send to repairing in micos

Day 3

Both spectral format and lt confirm that the grating is in the real blaze angle which is 75.22 and not the nominal 76
BTM tilt of 10 arcmin along vertical direction to center the FL vignetting (along main dispersion) with good result.
BTM tilt of 5 arcmin along horizontal direction to center the orders (note: attention that the beam has been raised so more critical the top edge of prism)
APSU counter clock of 1 deg to align the slits with the column of the CCD to compensate the new blaze angle
Mechanical metrology of the spectrograph as it is to confirm the position and acquisition of flat filed mech pos to be able to check the strange behavior
Best focus still to be identified
Definition of the steps to go in Vacuum in particular for the fibers with Ian

Comments are closed.